40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Aug 9 13:32:02 2010, nancy, Update, IOO, Mode Cleaner ASC 6x
    Reply  Mon Aug 9 15:54:43 2010, Koji, Update, IOO, Mode Cleaner ASC 
       Reply  Mon Aug 9 21:50:50 2010, nancy, Update, IOO, Mode Cleaner ASC 
Message ID: 3387     Entry time: Mon Aug 9 13:32:02 2010     Reply to this: 3388
Author: nancy 
Type: Update 
Category: IOO 
Subject: Mode Cleaner ASC 

 E-log entry for Friday - will attach more plots to this entry on wednesday after i am back   to the 40.

 
Started working at some 1030 hrs and recording the Open Loop Tfs for all 6 loops.
The control was not so good, and I lost the lock quite a number of times while measureing
WFS  did not converge when the spot was aligned to the center. But there was convergence to a non-center point. So if  the control system was switched on near those points, it was converging to that point.
 
Autolocker : switches WFS control on directly, whereas the best way is to gradually increase the gain to 1. Also, the autolocker code now needs to be changed to incoporate the switing off the MC2 oplev in down and switch it on in the up script.
 
After Koji locked the Reference Cavity in the evening, I resumed measurements for the Open Loop TFs.
 
Measurement of the Open Loop Transfer Functions :
 
 
noise waveform was generated using arbitrary wf generator and injected into each loop.
An LPF was applied to have max co-relation at minimum disturabnce. (thanks to Rana)
The Transfer functions, Co-relations and Power Spectra were then measured using the DTT.
 
 
Power Spectrum of the IN1, IN2 and EXC shows clearly the suppression of the noise, and OLTF shows the phase margins.
 
- Courtesy Rana again for suggesting the idea of plotting power spectra of all signals in the same graph.
 
Later in the night , Koji worked with me and we reflected upon all TFs and changed gains whereevr required according to the phase margin considerations from the Open Loop TFs.
We used the same output matrix given in the previous e-log.   
 
 
Final gains -
 
Alignment Gain in the WFS Master - 1.000
 
Loop Gain
WFS1 P 0.27
WFS1 Y 0.7
WFS2 P 0.15
WFS2 Y 0.110
MC 2OPLEV P -0.1
MC2 OPLEV Y -0.1
 
 
this measurement invloved locking the MC to the correct position, with the spot centered at both the WFS and the QPD. invloved some cheating (offsets) after we tried centering w/o offsets.
demod signal was also centered while alignment.
credits to Koji for getting the correct lock position and also staying with me till late night in the lab
 
Important Points to be noted
 
1. All loops' histories have to be cleared while swtiching them on.
2. turn the loop output before the loop input so that there is no remnant history in the loop.
2. Alignment gain was gradually increased to 1. and tehn the oplevs turned on.
 
 
Later measured teh PSD of  6 error signals under 3 conditions -
 
New Control ON
 
New Control OFF
 
Old Control ON
 
 
Also measured the time series for the MC_trans and MC_refl for the 3 conditions.
 
 
 Status MC_Trans  MC_REF 

  

New Control ON  trans_on.pdf refl_on.pdf   
New Control OFF trans_off.pdf  refl_off.pdf  
Old Control ON trans_old.pdf refl_old.pdf  
 
 
 
 
 
 
 
  

 

Attachment 5: refl_off.pdf  1.219 MB  | Show | Show all
ELOG V3.1.3-